OpenStudy (anonymous):

Copy Cats charges a $21 surcharge for a batch of colored copies plus $0.15 per page. Heritage Printing charges a $25.50 surcharge for a batch of colored copies plus $0.12 per page. How many pages of colored copies would make Heritage Printing cheaper than Copy Cats?

OpenStudy (anonymous):

Copy Cats = 21 + 0.15n Heritage Printing = 25.50 + 0.12n (Where 'n' is the number of pages) So: 25.50 + 0.12n < 21 + 0.15n Solve for n.

OpenStudy (anonymous):

cc = 21 + .15p hp = 25.5 +.12p 25.5 +.12p < 21 +.15p 4.5 < .03p 4.5/.03 <p 150 pages

OpenStudy (anonymous):

:-|

OpenStudy (schrodinger):

fewscrews, what?

OpenStudy (anonymous):

Solving an algebraic inequality is hardly a challenge for those who know how to do it. People should try to explain concepts and give suggestive nudges in the right directions.

OpenStudy (schrodinger):

That's understandable. I was just confused at the ambiguity of your response, lol. Regardless, I like the cooperative nature of the site if it's not used on test questions, and if it's put in that "nudge" frame of mind you suggested. Plus, it looks cool and it's a good way to just tinker with the idea of teaching people.